40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3201     Entry time: Mon Jul 12 22:01:13 2010
Author: Koji 
Type: Update 
Category: SUS 
Subject: One TT suspended. Still need fine alignment 

Jenne and Koji

We tweaked the alignment of the TT mirror.

First we put a G&H mirror, but the mirror was misaligned and touching the ECD as the magnet was too heavy. We tried to move the wires towards the magnet by 1mm.
It was not enough but once we moved the clamps towards the magnet, we got the range to adjust the pitching back and forth.
We tried to align it by the feaher touch to the clamp, we could not get close to the precision of 10mrad as the final tightening of the clamp screws did change the alignment.

We will try to adjust the fine alignment tomorrow again.

The damping in pitch, yaw and longitudinal looks quite good. We will also try to characterize the damping of the suspension using a simple oplev setup.

Attachment 1: IMG_2634.jpg  65 kB  Uploaded Mon Jul 12 23:26:13 2010  | Show | Show all
ELOG V3.1.3-